# Zhan Gaoa, Ji Hyung Leeb, Ziwei Meic, Zhentao Shic

### v2

@@ A.1 Proof of Lemma 1
A.2.1 Unit Root Regressors
@@
Lemma A.2.
Under Assumptions 14,
$\sup_{j\in\mathcal{M}_{x},t\in[n]}|x_{j,t-1}|$ $\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\sqrt{n\log p},$ (A.9)
$\sup_{j\in\mathcal{M}_{x},t\in[n]}\left|\sum_{t=1}^{n}x_{j,t-1}u_{t}\right|$ $\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}n(\log p)^{1+\frac{1}{2r}}.$ (A.10)
Furthermore, recall that $\widehat{\sigma}_{j}=\sqrt{n^{-1}\sum_{t=1}^{n}(w_{j,t-1}-\bar{w}_{j})^{2}}.$ Then
$n(\log p)^{-1}\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\inf_{j\in\mathcal{M}_{x}}\widehat{\sigma}_{j}^{2}\leq\sup_{j\in\mathcal{M}_{x}}\widehat{\sigma}_{j}^{2}\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}n(\log p),$ (A.11)
$\inf_{j\in\mathcal{M}_{z}}\widehat{\sigma}_{j}^{2}\stackrel{{\scriptstyle\mathrm{p}}}{{\asymp}}\sup_{j\in\mathcal{M}_{z}}\widehat{\sigma}_{j}^{2}\stackrel{{\scriptstyle\mathrm{p}}}{{\asymp}}1,$ (A.12)
$\sup_{j\in\mathcal{M}_{x},t\in[n]}\left|\sum_{t=1}^{n}\widehat{\sigma}_{j}^{-1}w_{j,t-1}u_{t}\right|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}(\log p)^{\frac{3}{2}+\frac{1}{2r}}.$ (A.13)
Proof of Lemma A.2.
Eq. (A.9) comes from MS24’s (B.12). (A.10) can be justified by MS24’s Propositions 1 and 3(b), and the latter also implies (A.13). (A.11) and (A.12) are deduced in the proof of MS24’s Theorem 3. ∎
@@
Proof of (A.14). Recall from (3.5) that $\rho_{\zeta}=1-C_{\zeta}/n^{\tau}$. When $j\in\mathcal{M}_{x}$, we have $\zeta{}_{j,t}=\sum_{s=1}^{t}\rho_{\zeta}^{s}e_{j,t-s}$ for $t\geq 1$, which is a partial sum of $\alpha$-mixing sup-exponential components $e_{j,t-s}$ weighted by $\rho_{\zeta}^{s}.$ Define
@@
Note that $\rho_{\zeta}^{s}e_{j,t-s}$ is sub-exponential with an exponentially decaying $\alpha$-mixing coefficient, and thus $\zeta{}_{j,t}$ is the partial sum of $t$ observations from a sub-exponential and $\alpha$-mixing time series. MS24’s Lemma B.2 yields
$\sup_{t\leq a_{n}}|\zeta{}_{j,t}|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\sqrt{a_{n}\cdot\log p}=O\left[n^{\tau/2}(\log p)^{3/2}\right].$ (A.18)
@@
$|\zeta{}_{j,t}|\leq\left|\sum_{s\leq a_{n}}\rho_{\zeta}^{s}e_{j,t-s}\right|+\left|\sum_{a_{n}<s\leq t}\rho_{\zeta}^{s}e_{j,t-s}\right|\leq\left|\sum_{s\leq a_{n}}\rho_{\zeta}^{s}e_{j,t-s}\right|+\rho_{\zeta}^{a_{n}}\left|\sum_{0<s\leq t-a_{n}}\rho_{\zeta}^{s-a_{n}}e_{j,t-s+a_{n}}\right|.$ (A.19)
@@
By (A.19), (A.20), and (A.21), it follows
$\sup_{a_{n}<t\leq n}|\zeta{}_{j,t}|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}n^{\tau/2}(\log p)^{3/2}$ (A.22)
and then (A.14) follows by (A.18) and (A.22).
@@
    for a fixed $j$, which has been established (A.14). With this result, we can deduce the same upper bound of $T_{1}$ in MS24’s Equation (B.13).
@@
    Change the expression of $T_{2}$ above into MS24’s Equation (B.12)
    $\sup_{k\in\mathcal{M}_{x},t\in[n]}\left|\sum_{t=G+1}^{n}e_{k,t}\sum_{r=t-G+1}^{t-1}\rho_{\zeta}^{t-1-r}e_{j,r}\right|.$
    The upper bound of the new expression of $T_{2}$ can be analyzed in the same way as MS24.
@@
Proof of (A.16). Note that when $j\in\mathcal{M}_{x}$, we have the following AR(1) representation
$\zeta_{j,t}=\rho_{\zeta}\zeta_{j,t-1}+e_{j,t}.$ (A.23)
@@
$x_{k,t-1}\zeta_{j,t}=\rho_{\zeta}x_{k,t-1}\zeta_{j,t-1}+x_{k,t-1}e_{j,t}$
where $\rho_{\zeta}$ is defined as (3.5). By $x_{k,t-1}\zeta_{j,t}=(x_{k,t}-e_{k,t})\zeta_{j,t},$ we obtain
$x_{k,t}\zeta_{j,t}=\rho_{\zeta}x_{k,t-1}\zeta_{j,t-1}+x_{k,t-1}e_{j,t}+e_{k,t}\zeta_{j,t}.$ (A.24)
Summing up both sides of (A.24) and using the fact that $\sum_{t=1}^{n}x_{k,t}\zeta_{j,t}=\sum_{t=1}^{n}x_{k,t-1}\zeta_{j,t-1}-x_{k,0}\zeta_{j,0}+x_{k,n}\zeta_{j,n}$, we deduce
$\sum_{t=1}^{n}x_{k,t-1}\zeta_{j,t-1}-x_{k,0}\zeta_{j,0}+x_{k,n}\zeta_{k,n}=\rho_{\zeta}\sum_{t=1}^{n}x_{k,t-1}\zeta_{j,t-1}+\sum_{t=1}^{n}x_{k,t-1}e_{j,t}+\sum_{t=1}^{n}e_{k,t}\zeta_{j,t}.$
@@
$(1-\rho_{\zeta})\sum_{t=1}^{n}x_{k,t-1}\zeta_{j,t-1}=\left(x_{k,0}\zeta_{j,0}-x_{k,n}\zeta_{j,n}\right)+\sum_{t=1}^{n}x_{k,t-1}e_{j,t}+\sum_{t=1}^{n}e_{k,t}\zeta_{j,t}.$ (A.25)
@@
$\sup_{k\in\mathcal{M}_{x}}\left|x_{k,0}\zeta_{j,0}-x_{k,n}\zeta_{j,n}\right|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}n^{\frac{1+\tau}{2}}(\log p)^{2}.$ (A.26)
@@
$\sup_{k\in\mathcal{M}_{x}}\left|\sum_{t=1}^{n}x_{k,t-1}e_{j,t}\right|+\sup_{k\in\mathcal{M}_{x}}\left|\sum_{t=1}^{n}e_{k,t}\zeta_{j,t}\right|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}n(\log p)^{1+\frac{1}{2r}}.$ (A.27)
@@
$(1-\rho_{\zeta})\sup_{k\in\mathcal{M}_{x}}\left|\sum_{t=1}^{n}x_{k,t-1}\zeta_{j,t-1}\right|$ $\leq\sup_{k\in\mathcal{M}_{x}}\left|x_{k,0}\zeta_{j,0}-x_{k,n}\zeta_{j,n}\right|+\sup_{k\in\mathcal{M}_{x}}\left|\sum_{t=1}^{n}x_{k,t-1}e_{j,t}\right|+\sup_{k\in\mathcal{M}_{x}}\left|\sum_{t=1}^{n}e_{k,t}\zeta_{j,t}\right|$
@@
Thus,
$\sup_{k\in\mathcal{M}_{x}}\left|\sum_{t=1}^{n}x_{k,t-1}\zeta_{j,t-1}\right|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\dfrac{n(\log p)^{1+\frac{1}{2r}}}{1-\rho_{\zeta}}\preccurlyeq n^{1+\tau}(\log p)^{1+\frac{1}{2r}}.$
given $1-\rho_{\zeta}=C_{\zeta}n^{-\tau}.$
Proof of (A.17). By the law of large number (LLN) in Phillips and Magdalinos, (2007, Eq. (4)), we have
$\dfrac{\sum_{t=1}^{n}\zeta_{j,t-1}^{2}}{n^{1+\tau}}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}\dfrac{{\rm var}(e_{j,t})}{2C_{\zeta}}.$ (A.28)
Without loss of generality, assume $\zeta_{j,0}=0$. Summing up both sides of (A.23), we have
$x_{j,n}-x_{j,0}=\sum_{t=1}^{n}e_{j,t}=\sum_{t=1}^{n}\zeta_{j,t}-\rho_{\zeta}\sum_{t=1}^{n}\zeta_{j,t-1}=\zeta_{j,n}+(1-\rho_{\zeta})\sum_{t=1}^{n}\zeta_{j,t-1}.$
Since $1-\rho_{\zeta}=C_{\zeta}n^{-\tau},$we have $\sum_{t=1}^{n}\zeta_{j,t-1}=n^{\tau}C_{\zeta}^{-1}(x_{j,n}-x_{j,0}-\zeta_{j,n}).$ By (A.9) and (A.14),
$\sum_{t=1}^{n}\zeta_{j,t-1}=O_{p}\left[n^{\tau}\cdot\left(\sqrt{n\log p}+n^{\tau/2}(\log p)^{3/2}\right)\right]=O_{p}\left(n^{\tau+1/2}\sqrt{\log p}\right),$ (A.29)
@@
$\dfrac{\widehat{\varsigma}_{j}^{2}}{n^{\tau}}=\dfrac{\sum_{t=1}^{n}\zeta_{j,t-1}^{2}}{n^{1+\tau}}-\dfrac{1}{n^{\tau}}\left(\dfrac{\sum_{t=1}^{n}\zeta_{j,t-1}}{n}\right)^{2}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}\dfrac{{\rm var}(e_{j,t})}{2C_{\zeta}}.$ (A.31)
@@
for an arbitrary $(p-1)$-dimensional vector $\varphi$. We can write the above inequality into
@@
$\|n^{-1}D_{-j}^{-1}W_{-j,\cdot}^{\top}\widetilde{\zeta}_{j}\|_{\infty}\leq(n\widehat{\varsigma}_{j})^{-1}\left(\sup_{k\in\mathcal{M}_{x}}\left|\sum_{t=1}^{n}\widehat{\sigma}_{k}^{-1}w_{k,t-1}\zeta_{j,t-1}\right|+\sup_{k\in[p_{z}]}\left|\sum_{t=1}^{n}\widehat{\sigma}_{k}^{-1}w_{k,t-1}\zeta_{j,t-1}\right|\right).$
Repeatedly using the bounds in (A.11), (A.12) in Lemma A.2, and (A.15), (A.16), and (A.17) in Lemma A.3,
@@
$\leq\left(\dfrac{\max_{j\in\mathcal{M}_{z},t\in[n]}\left|w_{j,t}\right|}{\inf_{j\in\mathcal{M}_{z}}\widehat{\sigma}_{j}}\vee\dfrac{\max_{j\in\mathcal{M}_{x},t\in[n]}\left|w_{j,t}\right|}{\inf_{j\in\mathcal{M}_{x}}\widehat{\sigma}_{j}}\right)\cdot\max_{k\in\mathcal{M}_{x},t\in[n]}\left|w_{k,t}\right|$
@@
where the fourth row applies (A.9), (A.11), (A.12), and (A.39). Following similar arguments we can deduce

### v3

@@ A.1.3 Proofs of Lemma 2
A.2.1 Local Unit Root Regressors
@@
Proof of (A.40). Recall from (3.5) that $\rho_{\zeta}=1-C_{\zeta}/n^{\tau}$. By (13) in phillips2009econometric, when $j\in\mathcal{M}_{x}$ we have
$\zeta{}_{j,t}=\zeta_{j,t}^{0}+\frac{C}{n}\psi_{j,t}^{0}\text{ for }t\geq 1,$ (A.44)
where
$\zeta_{j,t}^{0}=\sum_{s=1}^{t}\rho_{\zeta}^{s}e_{j,t-s}\text{ and }\psi_{j,t}^{0}=\sum_{s=0}^{t-1}\rho_{\zeta}^{s}x_{j,t-s-1}$
is a partial sum of $\alpha$-mixing sup-exponential components $e_{j,t-s}$ weighted by $\rho_{\zeta}^{s}.$
We first bound $\zeta_{j,t}^{0}$. Define
@@
Note that $\rho_{\zeta}^{s}e_{j,t-s}$ is sub-exponential with an exponentially decaying $\alpha$-mixing coefficient, and thus $\zeta^{0}{}_{j,t}$ is the partial sum of $t$ observations from a sub-exponential and $\alpha$-mixing time series. MS24’s Lemma B.2 yields
$\sup_{t\leq a_{n}}|\zeta_{j,t}^{0}|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\sqrt{a_{n}\cdot\log p}=O\left[n^{\tau/2}(\log p)^{3/2}\right].$ (A.45)
@@
$|\zeta_{j,t}^{0}|\leq\left|\sum_{s\leq a_{n}}\rho_{\zeta}^{s}e_{j,t-s}\right|+\left|\sum_{a_{n}<s\leq t}\rho_{\zeta}^{s}e_{j,t-s}\right|\leq\left|\sum_{s\leq a_{n}}\rho_{\zeta}^{s}e_{j,t-s}\right|+\rho_{\zeta}^{a_{n}}\left|\sum_{0<s\leq t-a_{n}}\rho_{\zeta}^{s-a_{n}}e_{j,t-s+a_{n}}\right|.$ (A.46)
@@
By (A.46), (A.47), and (A.48), it follows that
$\sup_{a_{n}<t\leq n}|\zeta_{j,t}^{0}|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}n^{\tau/2}(\log p)^{3/2}.$ (A.49)
By (A.45) and (A.49), we have
$\sup_{1<t\leq n}|\zeta_{j,t}^{0}|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}n^{\tau/2}(\log p)^{3/2}.$ (A.50)
We then bound $\psi_{j,t}^{0}$. By (A.25),
$\sup_{t\in[n]}\left|\psi_{j,t}^{0}\right|$ $\leq\sum_{s=0}^{n}\rho_{\zeta}^{s}\sup_{t\in[n]}\left|x_{j,t-s-1}\right|$
$\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\sqrt{n\log p}\cdot\sum_{s=0}^{n}\rho_{\zeta}^{s}$
$\leq\sqrt{n\log p}\cdot\frac{1}{1-\rho_{\zeta}}=O(n^{\frac{1}{2}+\tau}\sqrt{\log p}).$ (A.51)
Then by (A.44), (A.50) and (A.51),
$\sup_{t\in[n]}\left|\zeta_{j,t}\right|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}n^{\tau/2}(\log p)^{3/2}+n^{\tau-\frac{1}{2}}\sqrt{\log p}=O(n^{\tau/2}(\log p)^{3/2}),$
where the second step applies the fact that $\tau\in(0,1)$, which implies $\tau-\frac{1}{2}<\frac{\tau}{2}$. Then (A.40) follows.
@@
    for a fixed $j$, which has been established in (A.40). With this result, we can deduce the same upper bound of $T_{1}$ in MS24’s Equation (B.13).
@@
    Change the expression of $T_{2}$ above MS24’s Equation (B.12) into
    $T_{2}$ $=\sup_{k\in\mathcal{M}_{x},t\in[n]}\left|\sum_{t=G+1}^{n}e_{k,t}\sum_{r=t-G+1}^{t-1}\rho_{\zeta}^{t-1-r}\Delta x_{j,r}\right|$
    $\leq\sup_{k\in\mathcal{M}_{x},t\in[n]}\left|\sum_{t=G+1}^{n}e_{k,t}\sum_{r=t-G+1}^{t-1}\rho_{\zeta}^{t-1-r}e_{j,r}\right|+\dfrac{|c_{j}|}{n}\cdot\sup_{k\in\mathcal{M}_{x},t\in[n]}\left|\sum_{t=G+1}^{n}e_{k,t}\sum_{r=t-G+1}^{t-1}\rho_{\zeta}^{t-1-r}x_{j,r-1}\right|,$
    where the second step follows the DGP of LURs $x_{j,t}=(1-c_{j}^{*}/n)x_{j,t-1}+e_{j,t}$. Following the same way as MS24 to bound $T_{2}$, we can bound the first term by
    $\sup_{k\in\mathcal{M}_{x},t\in[n]}\left|\sum_{t=G+1}^{n}e_{k,t}\sum_{r=t-G+1}^{t-1}\rho_{\zeta}^{t-1-r}e_{j,r}\right|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}n\log p.$ (A.52)
    In addition, it is easy to show that
    $\sup_{k\in\mathcal{M}_{x},t\in[n]}|e_{k,t}|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\log p$ (A.53)
    given that $e_{k,t}$ is sub-exponential. Therefore, the second term is bounded by
    $n^{-1}\sup_{k\in\mathcal{M}_{x},t\in[n]}\left|\sum_{t=G+1}^{n}e_{k,t}\sum_{r=t-G+1}^{t-1}\rho_{\zeta}^{t-1-r}x_{j,r-1}\right|$ $\leq\sup_{k\in\mathcal{M}_{x}}n^{-1}\sum_{t=G+1}^{n}\left|e_{k,t}\right|\sup_{k\in\mathcal{M}_{x},t\in[n]}\sum_{r=t-G+1}^{t-1}|x_{j,r-1}|$
    $\leq G\cdot\sup_{k\in\mathcal{M}_{x},t\in[n]}|e_{k,t}|\cdot\sup_{t\in[n]}|x_{j,t}|$ (A.54)
    $\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\sqrt{n}(\log p)^{3/2},$ (A.55)
    where the last step follows (A.25) and (A.53). By (A.52) and (A.55), for our case we also have
    $T_{2}\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}n\log p.$
@@
Proof of (A.42). By the decomposition (A.44), we have
$\sup_{k\in\mathcal{M}_{x}}\left|\sum_{t=1}^{n}x_{k,t-1}\zeta_{j,t-1}\right|\leq\sup_{k\in\mathcal{M}_{x}}\left|\sum_{t=1}^{n}x_{k,t-1}\zeta_{j,t-1}^{0}\right|+\dfrac{C_{\zeta}}{n}\sup_{k\in\mathcal{M}_{x}}\left|\sum_{t=1}^{n}x_{k,t-1}\psi_{j,t-1}^{0}\right|.$ (A.56)
We first bound $\sup_{k\in\mathcal{M}_{x}}\left|\sum_{t=1}^{n}x_{k,t-1}\zeta_{j,t-1}^{0}\right|.$ Note that when $j\in\mathcal{M}_{x}$, we have the following AR(1) representation
$\zeta_{j,t}^{0}=\rho_{\zeta}\zeta_{j,t-1}^{0}+e_{j,t}.$ (A.57)
@@
$x_{k,t-1}\zeta_{j,t}^{0}=\rho_{\zeta}x_{k,t-1}\zeta_{j,t-1}^{0}+x_{k,t-1}e_{j,t}$
where $\rho_{\zeta}$ is defined in (3.5). By $x_{k,t-1}\zeta_{j,t}^{0}=(\rho_{k}^{*})^{-1}(x_{k,t}-e_{k,t})\zeta_{j,t}^{0},$ we obtain
$x_{k,t}\zeta_{j,t}^{0}=\rho_{k}^{*}\rho_{\zeta}x_{k,t-1}\zeta_{j,t-1}^{0}+\rho_{k}^{*}x_{k,t-1}e_{j,t}+e_{k,t}\zeta_{j,t}^{0}.$ (A.58)
Summing up both sides of (A.58) and using the fact that $\sum_{t=1}^{n}x_{k,t}\zeta_{j,t}^{0}=\sum_{t=1}^{n}x_{k,t-1}\zeta_{j,t-1}^{0}-x_{k,0}\zeta_{j,0}+x_{k,n}\zeta_{j,n}^{0}$, we deduce
$\sum_{t=1}^{n}x_{k,t-1}\zeta_{j,t-1}^{0}-x_{k,0}\zeta_{j,0}^{0}+x_{k,n}\zeta_{k,n}^{0}=\rho_{k}^{*}\rho_{\zeta}\sum_{t=1}^{n}x_{k,t-1}\zeta_{j,t-1}^{0}+\rho_{k}^{*}\sum_{t=1}^{n}x_{k,t-1}e_{j,t}+\sum_{t=1}^{n}e_{k,t}\zeta_{j,t}^{0}.$
@@
$(1-\rho_{k}^{*}\rho_{\zeta})\sum_{t=1}^{n}x_{k,t-1}\zeta_{j,t-1}^{0}=\left(x_{k,0}\zeta_{j,0}^{0}-x_{k,n}\zeta_{j,n}^{0}\right)+\rho_{k}^{*}\sum_{t=1}^{n}x_{k,t-1}e_{j,t}+\sum_{t=1}^{n}e_{k,t}\zeta_{j,t}^{0}.$ (A.59)
@@
$\sup_{k\in\mathcal{M}_{x}}\left|x_{k,0}\zeta_{j,0}^{0}-x_{k,n}\zeta_{j,n}^{0}\right|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}n^{\frac{1+\tau}{2}}(\log p)^{2}.$ (A.60)
@@
$\sup_{k\in\mathcal{M}_{x}}\left|\rho_{k}^{*}\sum_{t=1}^{n}x_{k,t-1}e_{j,t}\right|+\sup_{k\in\mathcal{M}_{x}}\left|\sum_{t=1}^{n}e_{k,t}\zeta_{j,t}^{0}\right|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}n(\log p)^{1+\frac{1}{2r}}.$ (A.61)
@@
$\sup_{k\in\mathcal{M}_{x}}\left|(1-\rho_{k}^{*}\rho_{\zeta})\sum_{t=1}^{n}x_{k,t-1}\zeta_{j,t-1}^{0}\right|$ $\leq\sup_{k\in\mathcal{M}_{x}}\left|x_{k,0}\zeta_{j,0}^{0}-x_{k,n}\zeta_{j,n}^{0}\right|+\sup_{k\in\mathcal{M}_{x}}\left|\sum_{t=1}^{n}x_{k,t-1}e_{j,t}\right|+\sup_{k\in\mathcal{M}_{x}}\left|\sum_{t=1}^{n}e_{k,t}\zeta_{j,t}^{0}\right|$
@@
Since $\sup_{k\in\mathcal{M}_{x}}\left|\frac{1}{1-\rho_{k}^{*}\rho_{\zeta}}\right|=O(n^{\tau})$, we have
$\sup_{k\in\mathcal{M}_{x}}\left|\sum_{t=1}^{n}x_{k,t-1}\zeta_{j,t-1}^{0}\right|\preccurlyeq n^{1+\tau}(\log p)^{1+\frac{1}{2r}}.$ (A.62)
In addition,
$\sup_{k\in\mathcal{M}_{x}}\left|\sum_{t=1}^{n}x_{k,t-1}\psi_{j,t-1}^{0}\right|$ $\leq n\sup_{k\in\mathcal{M}_{x},t\in[n]}|x_{k,t-1}|\cdot\sup_{k\in\mathcal{M}_{x},t\in[n]}|\psi_{j,t-1}^{0}|$
$\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}n\cdot\sqrt{n\log p}\cdot n^{\frac{1}{2}+\tau}\sqrt{\log p}=n^{2+\tau}\log p,$ (A.63)
where the second step applies (A.25) and (A.51). By (A.56), (A.62) and (A.63),
$\sup_{k\in\mathcal{M}_{x}}\left|\sum_{t=1}^{n}x_{k,t-1}\zeta_{j,t-1}\right|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}n^{1+\tau}(\log p)^{1+\frac{1}{2r}}+\frac{n^{2+\tau}\log p}{n}=O(n^{2+\tau}\log p).$
We complete the proof of (A.42).
Proof of (A.43). We have the following decomposition
$\dfrac{\widehat{\varsigma}_{j}^{2}}{n^{\tau}}=\dfrac{\sum_{t=1}^{n}\zeta_{j,t-1}^{2}}{n^{1+\tau}}-\dfrac{1}{n^{\tau}}\left(\dfrac{\sum_{t=1}^{n}\zeta_{j,t-1}}{n}\right)^{2}.$
By the law of large number (LLN) in phillips2009econometric, we have
$\dfrac{\sum_{t=1}^{n}(\zeta_{j,t-1})^{2}}{n^{1+\tau}}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}\dfrac{{\rm lvar}(e_{j,t})}{2C_{\zeta}},$ (A.64)
where ${\rm lvar}(e_{j,t})$ is the long-run variance of $e_{j,t}$.
We then bound $\sum_{t=1}^{n}\zeta_{j,t-1}.$ By (A.44), we have
$\sum_{t=1}^{n}\zeta_{j,t-1}=\sum_{t=1}^{n}\zeta_{j,t-1}^{0}+\frac{C_{\zeta}}{n}\sum_{t=1}^{n}\psi_{j,t-1}^{0}.$ (A.65)
We first bound $\sum_{t=1}^{n}\zeta_{j,t-1}^{0}$. Without loss of generality, assume $\zeta_{j,0}^{0}=0$. Summing up both sides of (A.57), we have
$\sum_{t=1}^{n}e_{j,t}=\sum_{t=1}^{n}\zeta_{j,t}^{0}-\rho_{\zeta}\sum_{t=1}^{n}\zeta_{j,t-1}^{0}=\zeta_{j,n}^{0}+(1-\rho_{\zeta})\sum_{t=1}^{n}\zeta_{j,t-1}^{0}.$
Since $1-\rho_{\zeta}=C_{\zeta}n^{-\tau},$we have $\sum_{t=1}^{n}\zeta_{j,t-1}^{0}=n^{\tau}C_{\zeta}^{-1}(\sum_{t=1}^{n}e_{j,t}-\zeta_{j,n}^{0}).$ Note that $\sum_{t=1}^{n}e_{j,t}$ is a unit root, and thus $\sum_{t=1}^{n}e_{j,t}=O_{p}(\sqrt{n\log p})$ by (A.25). Also, $|\zeta_{j,n}^{0}|=O_{p}(n^{\tau/2}(\log p)^{3/2})$ by (A.40). Therefore,
$\sum_{t=1}^{n}\zeta_{j,t-1}^{0}=O_{p}\left[n^{\tau}\cdot\left(\sqrt{n\log p}+n^{\tau/2}(\log p)^{3/2}\right)\right]=O_{p}\left(n^{\tau+1/2}\sqrt{\log p}\right).$ (A.66)
In addition, by (A.51)
$\sum_{t=1}^{n}\psi_{j,t-1}^{0}\leq n\cdot\sup_{t\in[n]}|\psi_{j,t-1}^{0}|=O_{p}(n^{3/2+\tau}\sqrt{\log p}).$ (A.67)
By (A.65), (A.66), and (A.67) we have
$\sum_{t=1}^{n}\zeta_{j,t-1}=O_{p}\left(n^{\tau+1/2}\sqrt{\log p}+\frac{n^{3/2+\tau}\sqrt{\log p}}{n}\right)=O_{p}\left(n^{\tau+1/2}\sqrt{\log p}\right),$ (A.68)
@@
$\dfrac{\widehat{\varsigma}_{j}^{2}}{n^{\tau}}=\dfrac{\sum_{t=1}^{n}\zeta_{j,t-1}^{2}}{n^{1+\tau}}-\dfrac{1}{n^{\tau}}\left(\dfrac{\sum_{t=1}^{n}\zeta_{j,t-1}}{n}\right)^{2}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}\dfrac{{\rm lvar}(e_{j,t})}{2C_{\zeta}}.$ (A.70)
@@
for an arbitrary (p1)(p-1)-dimensional vector $\varphi$. We can write the above inequality into
@@
$\|n^{-1}D_{-j}^{-1}W_{-j,\cdot}^{\top}\widetilde{\zeta}_{j}\|_{\infty}\leq(n\widehat{\varsigma}_{j})^{-1}\left(\sup_{k\in\mathcal{M}_{x}}\left|\sum_{t=1}^{n}\widehat{\sigma}_{k}^{-1}w_{k,t-1}\zeta_{j,t-1}\right|+\sup_{k\in\mathcal{M}_{z}}\left|\sum_{t=1}^{n}\widehat{\sigma}_{k}^{-1}w_{k,t-1}\zeta_{j,t-1}\right|\right).$
Repeatedly using the bounds in (A.35) and (A.36) in Lemma A.8, and (A.41), (A.42), and (A.43) in Lemma A.10,
@@
$\leq\left(\dfrac{\max_{j\in\mathcal{M}_{z},t\in[n]}\left|z_{j,t}\right|}{\inf_{j\in\mathcal{M}_{z}}\widehat{\sigma}_{j}}\vee\dfrac{\max_{j\in\mathcal{M}_{x},t\in[n]}\left|x_{j,t}\right|}{\inf_{j\in\mathcal{M}_{x}}\widehat{\sigma}_{j}}\right)\cdot\max_{k\in\mathcal{M}_{x},t\in[n]}\left|w_{k,t}\right|$
@@
where the last row applies (A.25), (A.35), (A.36), and (A.78). Following similar arguments we can deduce